### v1

@@ 6 Conclusion
In this study, we propose a two-stage ensemble modelling framework that addresses the challenge of incorporating multiple future covariates in spatio-temporal forecasting using INLA. Our approach integrates window-based XGBoost predictions as proxy covariates for future fire count and burnt area, and couples them with a latent Gaussian model to produce calibrated posterior forecasts. Furthermore, we introduce and implement the novel sub-asymptotic extended Generalised Pareto (eGP) likelihood within the INLA framework and its companion R-INLA library, enabling joint modelling of both moderate and extreme wildfire events.
By comparing posterior predictions under the eGP, Weibull, and Gamma likelihoods while keeping the remainder of the model structure fixed, we observe only marginal differences in predictive performance. We attribute this to the conditional independence assumption and the dominance of the linear predictor in shaping the marginal likelihoods. Similar findings are reported in Yadav et al., (2023), who also reported minimal sensitivity to likelihood choice in Bayesian hierarchical models for landslide size.
We also discuss a strategy for extending the framework to longer forecasting horizons by replicating the two-stage procedure for each horizon separately. While this does not offer a unified multi-horizon forecast, it provides a practical path forward within the current model constraints.
@@
7 Acknowledgment
RB work is partially funded by national funds through the FCT – Fundação para a Ciência e a Tecnologia, I.P., under the scope of the projects UIDB/00297/2020 (https://doi.org/10.54499/UIDB/00297/2020) and UIDP/00297/2020 (https://doi.org/10.54499/UIDP/00297/2020) (Center for Mathematics and Applications)
CC was partially funded by the Portuguese Fundação para a Ciência e a Tecnologia (FCT) I.P./MCTES through national funds (PIDDAC) – UID/50019/2025, LA/P/0068/2020 (https://doi.org/10.54499/LA/P/0068/2020)
@@ Appendix A Supplementary material
A.1 Code and Data
The code for implementing the two-stage model and reproducing the results in this paper is available at https://github.com/hcl516926907/Portugal_Wildfire. The wildfires and ERA5 data are publicly available online. Due to the size of the data, they are not shared in the above GitHub repository.
A.2 Derivation of the PC prior for $\kappa>0$
@@ A.3 Covariates in XGBoost
Pricp ERA5-Land $0.1^{\circ}\times 0.1^{\circ}$ Hourly Accumulated liquid and frozen water, including rain and snow, that falls to the Earth’s surface. Temp ERA5-Land $0.1^{\circ}\times 0.1^{\circ}$ Hourly Temperature of air at 2m above the surface of land. Ucomp ERA5-Land $0.1^{\circ}\times 0.1^{\circ}$ Hourly Eastward component of the 10m wind. Vcomp ERA5-Land $0.1^{\circ}\times 0.1^{\circ}$ Hourly Northward component of the 10m wind. DewPoint ERA5-Land $0.1^{\circ}\times 0.1^{\circ}$ Hourly Temperature to which the air, at 2 metres above the surface of the Earth, would have to be cooled for saturation to occur. HVegLAI ERA5-Land $0.1^{\circ}\times 0.1^{\circ}$ Hourly One-half of the total green leaf area per unit horizontal ground surface area for high vegetation type. LVegLAI ERA5-Land $0.1^{\circ}\times 0.1^{\circ}$ Hourly One-half of the total green leaf area per unit. horizontal ground surface area for low vegetation type. HVegCov ERA5 $0.25^{\circ}\times 0.25^{\circ}$ Constant The fraction of the grid box that is covered with vegetation that is classified as “high”. LVegCov ERA5 $0.25^{\circ}\times 0.25^{\circ}$ Constant The fraction of the grid box that is covered with vegetation that is classified as “low”. HVegTyp ERA5 $0.25^{\circ}\times 0.25^{\circ}$ Constant Indicator of the 6 types of high vegetation recognised by the ECMWF Integrated Forecasting System. LVegTyp ERA5 $0.25^{\circ}\times 0.25^{\circ}$ Constant Indicator of the 10 types of low vegetation recognised by the ECMWF Integrated Forecasting System. FWI NA $0.1^{\circ}\times 0.1^{\circ}$ Hourly Fire Weather Index RHumi NA $0.1^{\circ}\times 0.1^{\circ}$ Hourly Relative Humidity
Table A.2: Feature-engineered autoregressive covariates for fire count and burnt area. In this table, $t$ denotes the forecasting time point, $X$ indicates the spatial scale, taking values “dist” (district level) or “conc” (council level); and $Y$ specifies the source variable, with “fc” for fire count and “ba” for burnt area.
Name Index Range Formula Description X_Y_lag_j $j=1,2,\cdots,9$ $y_{s,t-j}$ Lag $j$ of monthly fire count/burnt area at council/district level X_Y_ma_j $j=3,6,9,12,24,36$ $\frac{\sum_{i=1}^{j}y_{s,t-i}}{j}$ Moving average of past $j$ months of fire count/burnt area at council/district level X_Y_hist_j $j=1,3,5$ $\frac{\sum_{k=1}^{3}\sum_{i=1}^{j}y_{s,t-12k+i-(j+1)/2}}{3j}$ Average fire count/burnt area over centred $j$ month around month $t$ in the past 3 years month_sin NA $\sin(2\pi t/12)$ Angular representation of the month month_cos NA $\cos(2\pi t/12)$ Angular representation of the month Year NA NA Year of the wildfire occurrence Lon NA NA Longitude of the centroid of the council Lat NA NA Latitude of the centroid of the council
A.4 Additional Diagnostic Plots

### v2

@@ 6 Conclusion
In this study, we propose a two-stage ensemble modelling framework that addresses the challenge of incorporating multiple future covariates in spatio-temporal forecasting using INLA. Our approach integrates window-based XGBoost predictions as proxy covariates for future fire count and burnt area, and couples them with a latent Gaussian model to produce calibrated posterior forecasts. By compressing complex spatio-temporal dynamics and dozens of environmental predictors into two informative proxy covariates, the framework enables INLA to exploit information that would otherwise be difficult to accommodate, thereby improving predictive performance. Furthermore, we introduce and implement the novel sub-asymptotic eGP likelihood within the INLA framework and its companion R-INLA library, enabling joint modelling of both moderate and extreme wildfire events.
By comparing posterior predictions under the eGP, Weibull, and Gamma likelihoods while keeping the remainder of the model structure fixed, we observe only a marginal improvement in predictive performance with the eGP likelihood. We attribute this to the conditional independence assumption and the dominance of the linear predictor in shaping the marginal likelihoods. Similar findings are reported in Yadav et al. (2023), who also reported minimal sensitivity to likelihood choice in Bayesian hierarchical models for landslide size.
We also discuss a strategy for extending the framework to longer forecasting horizons by replicating the two-stage procedure for each horizon separately. While this does not offer a unified multi-horizon forecast, it provides a practical path forward within the current model constraints. The result shows that our proposed framework can provide a stable forecast and uncertainty for longer horizons.
@@
7 Statements
7.1 Data Availability Statement
The code for implementing the two-stage model and reproducing the results in this paper is available at https://github.com/hcl516926907/Portugal_Wildfire. The wildfires data from ICNF (https://www.icnf.pt/florestas/gfr/gfrgestaoinformacao/estatisticas) and ERA5 data (https://cds.climate.copernicus.eu/datasets/reanalysis-era5-land?tab=overview) are publicly available online. Due to the size of the data, they are not shared in the above GitHub repository.
7.2 Funding Statement
  • RB work is partially funded by national funds through the FCT – Fundação para a Ciência e a Tecnologia, I.P., under the scope of the projects UIDB/00297/2020 (https://doi.org/10.54499/UIDB/00297/2020) and UIDP/00297/2020 (https://doi.org/10.54499/UIDP/00297/2020) (Centre for Mathematics and Applications)
  • CC was partially funded by the Portuguese Fundação para a Ciência e a Tecnologia (FCT) I.P./MCTES through national funds (PIDDAC) – UID/50019/2025, LA/P/0068/2020
7.3 Conflict of Interest Statement
We declare that they have no known competing financial interests or personal relationships that could have appeared to influence the work reported in this paper.
@@ Appendix A Supplementary material
A.1 Derivation of the PC prior for $\kappa>0$
@@ A.2 Covariates in XGBoost
Pricp ERA5-Land $0.1^{\circ}\times 0.1^{\circ}$ Hourly Accumulated liquid and frozen water, including rain and snow, that falls to the Earth’s surface. Temp ERA5-Land $0.1^{\circ}\times 0.1^{\circ}$ Hourly Temperature of air at 2m above the surface of land. Ucomp ERA5-Land $0.1^{\circ}\times 0.1^{\circ}$ Hourly Eastward component of the 10m wind. Vcomp ERA5-Land $0.1^{\circ}\times 0.1^{\circ}$ Hourly Northward component of the 10m wind. DewPoint ERA5-Land $0.1^{\circ}\times 0.1^{\circ}$ Hourly Temperature to which the air, at 2 metres above the surface of the Earth, would have to be cooled for saturation to occur. HVegLAI ERA5-Land $0.1^{\circ}\times 0.1^{\circ}$ Hourly One-half of the total green leaf area per unit horizontal ground surface area for high vegetation type. LVegLAI ERA5-Land $0.1^{\circ}\times 0.1^{\circ}$ Hourly One-half of the total green leaf area per unit. horizontal ground surface area for low vegetation type. HVegCov ERA5 $0.25^{\circ}\times 0.25^{\circ}$ Constant The fraction of the grid box that is covered with vegetation that is classified as “high”. LVegCov ERA5 $0.25^{\circ}\times 0.25^{\circ}$ Constant The fraction of the grid box that is covered with vegetation that is classified as “low”. HVegTyp ERA5 $0.25^{\circ}\times 0.25^{\circ}$ Constant Indicator of the 6 types of high vegetation recognised by the ECMWF Integrated Forecasting System. LVegTyp ERA5 $0.25^{\circ}\times 0.25^{\circ}$ Constant Indicator of the 10 types of low vegetation recognised by the ECMWF Integrated Forecasting System. FWI Derived $0.1^{\circ}\times 0.1^{\circ}$ Hourly Fire Weather Index RHumi Derived $0.1^{\circ}\times 0.1^{\circ}$ Hourly Relative Humidity
Table A.2: Feature-engineered autoregressive covariates for fire count and burnt area. In this table, $t$ denotes the forecasting time point, $h$ is the forecast horizon, $X$ indicates the spatial scale, taking values “dist” (district level) or “conc” (council level); and $Y$ specifies the source variable, with “fc” for fire count and “ba” for burnt area.
Name Index Range Formula Description X_Y_lag_j $j=1,2,\cdots,9$ $y_{s,t-j}$ Lag $j$ of monthly fire count/burnt area at council/district level X_Y_ma_j $j=3,6,9,12,24,36$ $\frac{\sum_{i=1}^{j}y_{s,t-i}}{j}$ Moving average of past $j$ months of fire count/burnt area at council/district level X_Y_hist_j $j=1,3,5$ $\frac{\sum_{k=1}^{3}\sum_{i=-(j-1)/2}^{(j-1)/2}y_{s,t+h-12k+i}}{3j}$ Average fire count/burnt area over centred $j$ month around month $t+h$ in the past 3 years month_sin NA $\sin(2\pi t/12)$ Angular representation of the month month_cos NA $\cos(2\pi t/12)$ Angular representation of the month Year NA NA Year of the wildfire occurrence Lon NA NA Longitude of the centroid of the council Lat NA NA Latitude of the centroid of the council
A.3 Additional Diagnostic Plots
@@
The uncertainty associated with the council-level spatio-temporal effect $G_{c}(\cdot)$ and the district-level spatio-temporal effect $G_{d}(\cdot)$ is illustrated below. Uncertainty is quantified using the 0.025 and 0.975 posterior quantiles of the corresponding random effects.
Figure A.2: Posterior uncertainty of the spatio-temporal effects. The top panel shows the 0.025 and 0.975 posterior quantiles of the council-level spatio-temporal effect $G_{c}(\cdot)$ grouped by month, while the bottom panel displays the averages of the 0.025 and 0.975 posterior quantiles of the district-level spatio-temporal effect $G_{d}(\cdot)$ aggregated over all time indices.